OpenStudy (anonymous):

Find the common difference for the given sequence. 1.05, 1.1, 1.15, 1.2, ...

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well the common difference is found by looking at the terms t2 - t1 t3 - t2 t4 - t3 1.10 - 1.05 = 1.15 - 1.10= 1.20 - 1.15 = do you get the same answer for each if so thats the common difference

OpenStudy (anonymous):

thanks campbell